OpenStudy (anonymous):

solve for x 2(x - 3)= -4x + 36

OpenStudy (anonymous):

2(x - 3)= -4x + 36 2x-6=-4x+36 2x+4x=36+6 6x=42 x=7

OpenStudy (anonymous):

2(x-3)=-4x+36 2x-6=-4x+36 +6 +6 2x=-4x+42 +4x +4x 6x=42 divide both sides by 6 x=7